! Required information For a bolted assembly with six bolts, the stiffness of each bolt is kb = 2.6 Mlbf/in and the stiffnessPLEASE PLEASE BE SURE THE ANSWER IS RIGHT, THANKS
Answer #1

GIVEN DATA

external load applied (p) = 80 kips

bolt stiffness ( Kb ) = 2.6(10^6) Ibf / in

Member stiffness (Km) = 15(10^6) Ibf / in

Diameter of bolts ( d ) = 1/2 in - 13 UNC grade 8

Number of bolts = 6

assumptions

for unified screw threads UNC and UNF

tensile stress area ( A ) = 0.1419 in^2

SAE specifications for steel bolts for grade 8

we have

Minimum proof strength ( Sp) = 120 kpsi

Minimum tensile strength (St) = 150 Kpsi

Load Bolt (p) = external load / number of bolts = 80 / 6 = 13.33 kips

Given the following values

Fi = 75%* Sp*At = (0.75*120*0.1419 ) = 12.771 kip

Preload stress

αi = 0.75Sp = 0.75 * 120 = 90 kpsi

stiffness constant

C = \frac{Kb}{Kb + Km}  = 2.6/17.6=0.147

A) yielding factor of safety

nP = \frac{sPAt}{Cp + Fi}

nP = 120*0.1419/((0.147*80)+12.771)=50.28/24.531=2.05Ans

Yielding factor of safety =2.05